If f(x) = \left\{\begin{matrix} x \left ( when 0 \leq x < 1 \right ) \\ f(x-1) \left ( when x\geq 1 \right ) \end{matrix}\right.

What is the value of f(4,7)? Explain and show your answers.
Answer #1

0 1 2 | flow - f(x-1) - زاده (ا لا ۰۰م 0.4 = لا - (0.1 + (۱۰ ع زار لا) 0.4 = (0.1) + = (ا- +++ = (1-1) زاده) - (204) : ((204